If you buy 6 pens and 1 mechanical pencil you will only get $1 change from your $10 dollar bill, but if you buy 4 pens and 2 mechanical pencils you will get $2 change. How much does each pen and mechanical pencil cost?
Answer by stanbon(75887)

If you buy 6 pens and 1 mechanical pencil you will only get $1 change from your $10 dollar bill, but if you buy 4 pens and 2 mechanical pencils you will get $2 change. How much does each pen and mechanical pencil cost?
------------------
Let # of pens be "p"; Let # of pencils be "c"
----------------------
EQUATIONS:
6p + c = 9
4p +2c = 8
------------
Rearrange:
6p+c = 9
2p+c = 4
----------
Substract 2nd from 1st to get:
4p = 5
p = 5/4 = $1.25 (price of a pen)
-----------
Substitute to solve for "c":
4(5/4)+2c = 8
5 + 2c = 8
2c = 3
c = 3/2 = $1.50 (price of a pencil)
============
Cheers,
Stan H.